OpenStudy (anonymous):

Help! Which of these is equivalent to log(3y)*2x a.) log(2x)/log(3y) b.) log(3y)/log(2x) c.) log*2x/3y d.) log*3y/2x

OpenStudy (anonymous):

Also how would you find out?

OpenStudy (lgbasallote):

do you mean \[\LARGE \log_{3y} 2x\]

OpenStudy (anonymous):

yes, how would you usually write it?

OpenStudy (lgbasallote):

it's written log_(3y) (2x)

OpenStudy (anonymous):

ohh okay

OpenStudy (lgbasallote):

anyway...you use change-of-base formula \[\huge \log_a b = \frac{\log_c b}{\log_c a}\] for example you have \[\large \log_x 2\] i can rewrite this as\[\Large \frac{\log_y 2}{\log_y x}\] or simply \[\Large \frac{\log 2}{\log x}\] does that give you any ideas?

OpenStudy (anonymous):

so its a!

OpenStudy (lgbasallote):

YES!!!

OpenStudy (anonymous):

yaay thank you!

OpenStudy (lgbasallote):

you're welcome ^_^

OpenStudy (anonymous):

wait I have one more question and I think it's basically the same

OpenStudy (anonymous):

Which of the following is equivalent to log_(7) (50) rounded to three decimal places? a.) 2.010 b.) 1.772 c.) 0.854 d.) 0.497

OpenStudy (anonymous):

do you do it the same way because in that case it would be 7.14 but those aren't one of the options

OpenStudy (lgbasallote):

\[\log_7 (50) = \frac{\log 50}{\log 7} \implies \log 50 \div \log 7\] is that what you did?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

isnt that what youre supposed to do

OpenStudy (lgbasallote):

actually... i think what you did was \(50 \div 7\)

OpenStudy (lgbasallote):

recheck what you did in your calculator

OpenStudy (anonymous):

ohhhhhh you have to put the log sign in?

OpenStudy (lgbasallote):

lol of course =)) \(\log 50 \div \log 7\) exactly that

OpenStudy (anonymous):

ya its a

OpenStudy (anonymous):

thank you for your help but will you maybe help me with one more thing?

OpenStudy (lgbasallote):

i'll try

OpenStudy (anonymous):

To find out how long it will take for $500 to double when invested at 5% annual interest compounded twice a year, you solve the following equation: 1000 = 500 (1+(.05/2))^2t. Use complete sentence to describe each step of math needed to solve this equation.

OpenStudy (anonymous):

I already did it but I feel like i did it really wrong..

OpenStudy (lgbasallote):

ugh sorry im not good with "compounded" thingies

OpenStudy (lgbasallote):

i dont know that

OpenStudy (anonymous):

okay no probs thankyou!